The normal boiling point of cyclohexane is 80.74 °C.
At 20.00 °C, the vapour pressure of cyclohexane is 0.1026 atm.
What is the vapour pressure of cyclohexane when the temperature is 30.00 °C?
Answer #1

We have to solve this problem by using integrated form of Clausius-Clapeyron equation....

Given, At, T = 20°C . P = 0.1026 atue l = 0.1026 atu 220 +273 l 293k T = 30.00c ,Py = ? = (30+273)K = 303k And, at normal boor 2.276 = 1H X 0.000586 8.314] mol or AH = 2.276x8.314 I mot 0.000586 . : AH = 32291.23 I moll Now how turn value, E T=2934,